In March, Kelly Company had the following unit production costs: materials $10 and conversion costs $9. On March 1, it had no work in process. During March, Kelly transferred out 23,000 units. As of March 31, 4,500 units that were 33% complete as to conversion costs and 100% complete as to materials were in ending work in process.

Compute the total units to be accounted for.

Total units Type your answer here

eTextbook and Media

  

  

Compute the equivalent units of production.

Materials

Conversion costs

The equivalent units of production

eTextbook and Media

  

  

Prepare a cost reconciliation schedule, including the costs of materials transferred out and the costs of materials in process.

Cost Reconciliation Schedule

Costs accounted for

   Transferred out

$

   Work in process, March 31

      Materials

$

      Conversion costs

Total costs

$
Answer #1

1) Total units to be accounted for =

23000+4500=27500 units

2) Equivalent unit of production

Total material equivalent units = finished units +100% work in process units

= 23000+4500=27500 units

Total conversion equivalent units = finished units +33% work in process units

= 23000+33%*4500

= 23000+1485 =24485 units

3) Cost reconcilliation schedule

  • Cost accounted for transferred out = units*total cost of material and conversion

= 23000 units *19

=$4,37,000

  • Work in process , march 31

material =4500*10 = $45000

conversion cost = 1485*9 = $13365

Cost Reconcilliation schedule

cost accounted for transferred out =$437000

+ work in process , march 31 =$58365

material $45000

conversion $13365

TOTAL COST = $495365

If you find answer suitable please provide feedback